OpenStudy (anonymous):

Find the area under P=90(0.7)^t between t=0 and t=6. please help!Medals are awarded!

OpenStudy (anonymous):

90*(0.7)^6/ln0.7

OpenStudy (anonymous):

Area under a curve can be found my integrating that curve. Set up your integral with the limits given, 0 and 6. \[\int\limits_{0}^{6}90(0.7)^tdt = 90\int\limits_{0}^{6}(0.7)^tdt\]

OpenStudy (anonymous):

\[90\int\limits_0^6 0.7^t \, dt\]

OpenStudy (anonymous):

why is it natural log and the other post 900 though?

OpenStudy (anonymous):

900 was wrong. A mistype.

OpenStudy (anonymous):

I meant 90 0.7 but space did not appear

OpenStudy (anonymous):

okay:) but why is it natural log?

OpenStudy (anonymous):

\[\int 0.7^tdt = \int e^{t \log{0.7}} = \frac{e^{t \log{0.7}}}{\log{0.7}} = \frac{0.7^t}{\log{0.7}}.\]

OpenStudy (anonymous):

hello every one i have given the correct ans...every1 is wrong